emmacole0307:

The total mass of the hydrogen gas in the Hindenburg zeppelin was 18,000 kg. What volume did the hydrogen occupy? (Assume that the temperature of the hydrogen was 0°C and that it was at a pressure of 1 atm.)

Narad:

The conditions are STP (0ºC and 1atm.) The molar volume is V= 22.4L Mass of hydrogen is = \[18*10^{6}\] g mole of hydrogen = 2g Number of moles of hydrogen =n= \[9* 10^{6}\] moles Volume of hydrogen = V*n =\[201.6* 10^{6}\] L =\[201.6 *10^{3} m ^{3}\]